ID |
Date |
Author |
Type |
Category |
Subject |
12053
|
Tue Mar 29 03:16:21 2016 |
ericq | Update | LSC | DRFPMI Locked Once More |
[ericq, Gautam]
Three RF-only locks longer than a minute tonight, out of 5 total attempts.
Last week, I determined that the beam spot on the RF POP PD is too large. This still needs to be fixed. I updated the ASS model to use REFLDC as a PRCL dither error signal; it works.
There seems to be some excess angular motion of ETMY tonight. This is evident in the oplev spectra (as compared to ETMX), and the GTRY camera, and even the retroreflected beam from a misalgined ETMY on the ITMY face when the PRC is carrier locked.
Gautam and I mostly focused on setting up the CAL-DARM_CINV block to produce this (mostly) calibrated spectrum starting from GPS 1143274087. [Darm on unwhitened AS55, DRMI on 3F, one CARM boost]

Here are the control and error signal spectra:

[DTT files attached]
Note to self: archive some of this data |
Attachment 1: 2016-03-29_calibdarm.pdf
|
|
Attachment 2: 2016-03-29_DRFPMI_errctrl.pdf
|
|
Attachment 3: DRFPMI_DTT.zip
|
12052
|
Mon Mar 28 22:16:44 2016 |
Koji | Update | General | New WiFi router |
I configured three more mini wifi extender. They are ready to use.
We should add these to the host table (I forgot where it is)
192.168.113.233 NETGEAR_EX3700_1
192.168.113.234 NETGEAR_EX3700_2
192.168.113.235 NETGEAR_EX3700_3
192.168.113.236 NETGEAR_EX3700_4
|
12051
|
Mon Mar 28 10:43:18 2016 |
Steve | Update | endtable upgrade | ETMX 4'x2' optical table pictures |
|
Attachment 1: ETMX4x2Layout.jpg
|
|
Attachment 2: ETMX_4x2.JPG
|
|
Attachment 3: ETMXopt_4x2.JPG
|
|
12050
|
Mon Mar 28 08:30:09 2016 |
Steve | Update | SUS | PRM damping restored |
Recent EQ 4.8 mag San Felipe, Mexico trips PRM sus damping.
PRM damping restored. PMC locked. |
Attachment 1: shaking.png
|
|
12049
|
Sat Mar 26 18:28:24 2016 |
Koji | Update | elog | elogd flakiness |
Elogd have been restarted several times today because it died everytime I submit something.
Here is the copy of the log.
GET /OMC_Lab/255?cmd=loc&value=Submit HTTP/1.1
Returned 6 bytes
GET /40m/elog.rdf HTTP/1.1
Returned 17109 bytes
TCP connection #1 on socket 5 closed
POST /OMC_Lab/ HTTP/1.1
Returned 20 bytes
GET /OMC_Lab/255 HTTP/1.1
Returned 53721 bytes
GET /ckeditor/skins/moono/images/arrow.png HTTP/1.1
Returned 489 bytes
POST /OMC_Lab/ HTTP/1.1
*** buffer overflow detected ***: /export/home/elog/elog/elogd terminated
======= Backtrace: =========
/lib/x86_64-linux-gnu/libc.so.6(__fortify_fail+0x37)[0x7f1435639e57]
/lib/x86_64-linux-gnu/libc.so.6(+0x108d50)[0x7f1435638d50]
/lib/x86_64-linux-gnu/libc.so.6(+0x1081b9)[0x7f14356381b9]
/lib/x86_64-linux-gnu/libc.so.6(_IO_default_xsputn+0xdd)[0x7f14355ab0cd]
/lib/x86_64-linux-gnu/libc.so.6(_IO_vfprintf+0x25a8)[0x7f143557ac18]
/lib/x86_64-linux-gnu/libc.so.6(__vsprintf_chk+0x94)[0x7f1435638254]
/lib/x86_64-linux-gnu/libc.so.6(__sprintf_chk+0x7d)[0x7f143563819d]
/export/home/elog/elog/elogd[0x426405]
/export/home/elog/elog/elogd[0x473b7f]
/export/home/elog/elog/elogd[0x4abfb2]
/export/home/elog/elog/elogd[0x4ad7fb]
/export/home/elog/elog/elogd[0x4b0af5]
/export/home/elog/elog/elogd[0x4b1eb9]
/export/home/elog/elog/elogd[0x403568]
/lib/x86_64-linux-gnu/libc.so.6(__libc_start_main+0xed)[0x7f143555176d]
/export/home/elog/elog/elogd[0x404299]
======= Memory map: ========
00400000-004e6000 r-xp 00000000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
006e5000-006e6000 r--p 000e5000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
006e6000-007c6000 rw-p 000e6000 fc:00 10361276 /export/home/elog/elog-3.0.d/elogd
007c6000-0173d000 rw-p 00000000 00:00 0
0214d000-02656000 rw-p 00000000 00:00 0 [heap]
7f14342f8000-7f143430d000 r-xp 00000000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143430d000-7f143450c000 ---p 00015000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450c000-7f143450d000 r--p 00014000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450d000-7f143450e000 rw-p 00015000 fc:00 2883628 /lib/x86_64-linux-gnu/libgcc_s.so.1
7f143450e000-7f14348cd000 rw-p 00000000 00:00 0
7f1434a34000-7f1434d39000 r--p 00000000 fc:00 530477 /usr/lib/locale/locale-archive
7f1434d39000-7f1434d4f000 r-xp 00000000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434d4f000-7f1434f4e000 ---p 00016000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f4e000-7f1434f4f000 r--p 00015000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f4f000-7f1434f50000 rw-p 00016000 fc:00 655527 /usr/local/lib/libz.so.1.2.8
7f1434f50000-7f1434f52000 r-xp 00000000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1434f52000-7f1435152000 ---p 00002000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435152000-7f1435153000 r--p 00002000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435153000-7f1435154000 rw-p 00003000 fc:00 2883655 /lib/x86_64-linux-gnu/libdl-2.15.so
7f1435154000-7f1435307000 r-xp 00000000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435307000-7f1435506000 ---p 001b3000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435506000-7f1435521000 r--p 001b2000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f1435521000-7f143552c000 rw-p 001cd000 fc:00 2883609 /lib/x86_64-linux-gnu/libcrypto.so.1.0.0
7f143552c000-7f1435530000 rw-p 00000000 00:00 0
7f1435530000-7f14356e4000 r-xp 00000000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14356e4000-7f14358e3000 ---p 001b4000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e3000-7f14358e7000 r--p 001b3000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e7000-7f14358e9000 rw-p 001b7000 fc:00 2884139 /lib/x86_64-linux-gnu/libc-2.15.so
7f14358e9000-7f14358ee000 rw-p 00000000 00:00 0
7f14358ee000-7f1435943000 r-xp 00000000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435943000-7f1435b42000 ---p 00055000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b42000-7f1435b45000 r--p 00054000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b45000-7f1435b4c000 rw-p 00057000 fc:00 2884155 /lib/x86_64-linux-gnu/libssl.so.1.0.0
7f1435b4c000-7f1435b6e000 r-xp 00000000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d57000-7f1435d5c000 rw-p 00000000 00:00 0
7f1435d6a000-7f1435d6e000 rw-p 00000000 00:00 0
7f1435d6e000-7f1435d6f000 r--p 00022000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7f1435d6f000-7f1435d71000 rw-p 00023000 fc:00 2884145 /lib/x86_64-linux-gnu/ld-2.15.so
7ffd85795000-7ffd85997000 rw-p 00000000 00:00 0 [stack]
7ffd859b2000-7ffd859b4000 r-xp 00000000 00:00 0 [vdso]
ffffffffff600000-ffffffffff601000 r-xp 00000000 00:00 0 [vsyscall]
er_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "cache_disable"
Received unknown cookie "NO_CACHE"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id"
Received unknown cookie "__utma"
Received unknown cookie "_ga"
Received unknown cookie "__unam"
Received unknown cookie "__utma"
Received unknown cookie "tk_ni"
Received unknown cookie "ajs_user_id"
Received unknown cookie "ajs_group_id"
Received unknown cookie "ajs_anonymous_id" |
12048
|
Fri Mar 25 23:54:04 2016 |
rana | Update | PEM | Guralp Seismometers |
Something seems not right. The Guralp response should be flat in velocity from 0.05-30 Hz. Why is there any feature at 1 Hz? Saturation of some kind? |
12047
|
Fri Mar 25 19:17:28 2016 |
Nikhil | Update | PEM | Guralp Seismometers |
Calibration of Guralp Seismometers
Objective
- Estimate transfer functions of Guralp A ( near ETMX) and Guralp B ( near ETMY)
- Calibrate the instruments by estimating Velocity Sensitity Parameter
- Convert previously measured Voltage Spectrum to Velocity Spectrum
Instruments Used
- Guralp CMG-40 T Seimometers : Guralp A (Serial Number: T4Q17)
- Guralp CMG-40 T Seimometers : Guralp B (Serial Number: T4157)
- Guralp Handheld Control Unit (HCU)
- FFT Spectrum Analyzer: Model SR785: 2 Channel Dynamic Signal Analyzer
- Oscilloscope: TDS 3014B
- Function Generator: DS 345
Procedure & Results
Sinusoidal current of known frequency and amplitude was injected to the Seismometer calibration coil using signal generator and handheld control unit & corresponding Magnitude and Phase response were recorded. For Guralp B, system response was also estimated with a FFT Spectrum Analyzer.

Frequnecy Range: 0.1 Hz to 45 Hz.
Equivalent Input Velocity was derived from the Input Voltage measurements using the relation: v = V/ (2*pi*f*R*K) , V is the peak to peak Calibration Signal voltage, f is the calibration signal frequency, R is the calibration resistor and K is the feedback coil constant. [See Appendix for R & K values]
Velocity Sensitity at the required frequency is obtained by dividing the Output Response Voltage by the Equivalent Input Velocity.

The obtained Velocity Sensitivity is used to convert the recorded Volatge PSD to Velocity PSD as shown below. The obtained results are compared to gloabl high noise model [NHNM] and USGS New Low Noise Model [NLNM,Peterson 1993] which gives the lowest observed vertical seismic noise levels across the seismic frequency band. Plot legend NLNM shows both the high & low levels.
Guralp A [X Arm] Low Velocity Output

Guralp B [Y Arm] Low Velocity Output

DTT Power Spectrum
Both the Seismometers were connected to the 40 M Control and Data Acquisition System (CDS) and Power Spectrum was estimated for the Vertical, North/South & East/West Channels using Diagnostic Test Tool (DTT) software.

Comments
- The transfer function from Guralp A [ETMX] looks similar to that of Guralp B [ETMY] in both magnitude and phase but with a lower gain.
- Velocity Sensitivity of Guralp A is comparable to the value provided in the Calibration Data Sheet [~ 400] for all the channels [Vertical, North/South, East/West] after 1 Hz. For Guralp B, Velocity Sensitivity is a factor of 2.5 higher [all channels] than the specification [~ 400] after 1 HZ.Below 1 Hz Sensitivity drops down for both sensors. I am not ruling out a missing common factor in the calculation, but anyway, test shows that Guralp B has ~2.5 times better Velocity Sensitivity than Guralp A.
- The Calibrated Seismic Velocity Spectrum for Guralp B is within the Globally Observed High and Low Noise Seismic Spectrum while Guralp A's Spectrum is more noisier above 1 Hz [Anthropogenic Activity normally contributes the most in 1 Hz to 10 Hz frequency band].
- Concurrently acquired Power Spectrum using DTT [Diagnostic Test Tools] shows that Guralp A Spectrum behaves rather strangely. The system response seems to be completely different from the one we obtained locally using signal generator. While Guralp B functionality seems normal. One reason for this erratic beahvior might be faulty cables used for data acquisition from Guralp A. This needs to be verified.
Appendix
CMG-40T Guralp A Calibration Sheet
|
Velocity Output: V/m/s (Differential) |
Mass Position Output (Acceleration Output) : V/m/s^2 |
Feedback Coil Constant : Amp/m/s^2 |
Vertical |
2 x 400 |
19 |
0.00397 |
North/South |
2 x 398 |
23 |
0.00486 |
East/West |
2 x 401 |
23 |
0.00479 |
Calibration Resistor: 51000
CMG-40T Guralp B Calibration Sheet
|
Velocity Output: V/m/s (Differential) |
Mass Position Output (Acceleration Output) : V/m/s^2 |
Feedback Coil Constant : Amp/m/s^2 |
Vertical |
2 x 401 |
19 |
0.00408 |
North/South |
2 x 400 |
20 |
0.00421 |
East/West |
2 x 404 |
22 |
0.00466
|
Calibration Resistor: 51000 |
Attachment 3: VelSens_XArm_Guralp_A.png
|
|
Attachment 4: VelSens_XArm_Guralp_A.png
|
|
Attachment 5: VelSens_YArm_Guralp_B.png
|
|
Attachment 8: Vel_PSD_XArm_GurB_E.png
|
|
Attachment 9: Vel_PSD_XArm_GurB_N.png
|
|
Attachment 16: Guralp_PowerSpectrum.pdf
|
|
12046
|
Thu Mar 24 08:20:52 2016 |
Steve | Update | PEM | Guralp-A calibration sheet |
All Guralp instruments and digitisers are provided with calibration documentation. Should you require a copy of calibration information for any product, email caldoc@guralp.com with the serial number of the product in the subject field and calibration information will be sent to you through email.
See data in the 40m wiki
|
12045
|
Thu Mar 24 07:56:09 2016 |
Steve | Update | Calibration-Repair | NO Noise Eater for 1W Innolight |
1W Innolight is NOT getting Noise Eater as it was decided yesterday at the 40m meeting. Corrected 3-25-2016
Repair quote with adding noise eater is in 40m wiki
Quote: |
Quote: |
Quote: |
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I
|
Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010
It's diodes should be replaced, based on it's age and performance.
RIN and noise eater bad. I will get a quote on this job.
The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956
|
Diagnoses from Glasglow:
“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.”
|
|
12044
|
Wed Mar 23 15:23:12 2016 |
Steve | Summary | PEM | Guralps as connected |
We have one calibration sheet of GUR- B, from 26 June 2008, model CMG-T40-0008, sn T4157 at ETMY east, interface box input 1
I'm looking for calibration paper of GUR- A, model CMG-T40-0053, sn T4Q17 at ETMX south, interface box input 2
Quote: |
I measured the guralp raw outputs and the TFs using the handheld unit and an FFT analyzer.
[Setup]
The handheld unit was connected to each guralp with the same cable which is confirmed t be functional with the Yend Guralp.
The signal for Z, N, and E directions are obtained from the banana connectors on the handheld unit. Each direction has mass, low gain velocity, and high gain velocity output. The PSDs of the signals were measured with an FFT analyzer. The transfer function from the mass signal to the low/high gain signals were also measured for each direction.
The adjustment screw for the E output of the Xend does not work. I had to tilt the Xend Guralp using the leg screws to bring the E signal to zero.
[Result]
Attachment 1: Raw voltage PSD for all outputs
Attachment 2: Comparison of the low gain vel outputs
- All of the mass output show similar PSDs.
- Low gain velocity outputs shows somewhat similar levels. I still need to check if the output is really the ground velocity or not.
- High gain velocity outputs are either not high gain, broken, or not implemented.
- We need to calibrate the low gain output using signal injection, huddle test, or something else.
Attachment 3: TFs between each mass output and the low or high gain outputs
- TFs between the mass signal and the low vel signals show the similar transfer functions between the channels.
- The high gain outputs show low or no transfer function with regard to the mass signals.
|
|
Attachment 1: GUR_A.jpg
|
|
Attachment 2: GUR_B.jpg
|
|
Attachment 3: GUR_Interface_Box.jpg
|
|
12043
|
Wed Mar 23 11:55:47 2016 |
Steve | Update | General | Smart UPS 2200 Battery Replaced |
Batteries replaced in control room UPS after 3 years from replaceUPSbattery.com
|
12042
|
Tue Mar 22 21:30:15 2016 |
Koji | Update | IOO | PMCIMC aligned, WFS offset adjusted |
The alignment of the PMC adjusted on the PSL table: Trans 0.737->0.749
The alignment of the IMC adjusrted on the sliders: Trans 14300->15300
WFS offset has been reset by /opt/rtcds/caltech/c1/scripts/MC/WFS/WFSoffsets |
Attachment 1: 08.png
|
|
12041
|
Tue Mar 22 14:12:18 2016 |
Steve | Update | IOO | Laser is turned back on |
The 2W Innolight was turned on.
|
Attachment 1: off-onAgain.png
|
|
12040
|
Mon Mar 21 14:29:32 2016 |
Steve | Update | Calibration-Repair | 1W Innolight laser repair diagnoses |
Quote: |
Quote: |
After adjusting the alignment of the two beams onto the PD, I managed to recover a stronger beatnote of ~ -10dBm. I managed to take some measurements with the PLL locked, and will put up a more detailed post later in the evening. I turned the IMC autolocker off, turned the 11MHz Marconi output off, and closed the PSL shutter for the duration of my work, but have reverted these to their nominal state now. The are a few extra cables running from the PSL table to the area near the IOO rack where I was doing the measurements from, I've left these as is for now in case I need to take some more data later in the evening...I
|
Innolight 1W 1064nm, sn 1634 was purchased in 9-18-2006 at CIT. It came to the 40m around 2010
It's diodes should be replaced, based on it's age and performance.
RIN and noise eater bad. I will get a quote on this job.
The Innolight Manual frequency noise plot is the same as Lightwave' elog 11956
|
Diagnoses from Glasglow:
“So far we have analyzed the laser. The pump diode is degraded. Next we would replace it with a new diode. We would realign the diode output beam into the laser crystal. We check all the relevant laser parameters over the whole tuning range. Parameters include single direction operation of the ring resonator, single frequency operation, beam profile and others. If one of them is out of spec, then we would take actions accordingly. We would also monitor the output power stability over one night. Then we repackage and ship the laser.” |
12039
|
Mon Mar 21 10:07:45 2016 |
Steve | Update | SUS | ETMX enclosure update |
The enclosure top piece in the middle is still in the machine shop.
The carpenter who helps in the built just left for one week vacation.
The unit will be ready on April 1
I'd prefer doing the installation with the enclosure on the new table.
It's the only way to minimize the resonances of the enclosure with shimming. |
Attachment 1: ETMXenclosure.jpg
|
|
12038
|
Fri Mar 18 11:04:56 2016 |
Steve | Update | SUS | inverted pendulum |
Kate Dooley picked up this item today. |
Attachment 1: invPeng.jpg
|
|
12037
|
Wed Mar 16 16:02:40 2016 |
Koji | Update | SUS | wire standoff test cut |
It looks almost OK, but we need a bit sharper picture for both the groove and thw wire. |
Attachment 1: ruby.png
|
|
12036
|
Wed Mar 16 15:36:03 2016 |
Steve | Update | SUS | wire standoff test cut |
Ruby wire standoff 1 mm od. with V-groove test cut. SOS sus wire 0.0017" od. is in the background.
|
Attachment 1: 0182_-_20160316_135935.png
|
|
12035
|
Tue Mar 15 10:31:58 2016 |
Steve | Update | IOO | Laser is turned back on |
It's may be the janitor's doing.
I noticed that the HEPA filers were off. They are turned on at 20%
|
Attachment 1: 2WlaserOff-On.png
|
|
12033
|
Mon Mar 14 22:42:23 2016 |
gautam | Update | endtable upgrade | Inventory check |
Quote: |
Steve should be able to get another copy of the EY doubler mount made up if we really don't have another one sitting in the Manasa end table box which Koji mentioned.
|
I located the second doubler mount, it was sitting inside a cabinet along the Y-arm. So this will not have to be machined. The doubling oven mount is black in colour.
So as things stand now, the only thing that needs to be machined is a non-green mount for the IR faraday (IO-5-1064-HP) - is it possible to just coat the existing mount with a different color? I've got a drawing for this part ready, but it seems unnecessary to machine the whole thing from scratch when only the color is an issue. Steve was talking about dipping this in some sort of solution and taking the green off. But if this isn't possible, I'll send Steve the drawings tomorrow so that he can place the order with the machine shop...
I will work on the mode-matching calculations over the next couple of days to make sure we have all the mirrors and lenses we need.
|
12032
|
Sat Mar 12 22:23:37 2016 |
rana | Summary | IOO | PMC relocked |
Found it locked on TEM01 mode.
Sweets in the fridge for non-PhD holders, courtesy of the highest levels of Caltech. |
12031
|
Fri Mar 11 16:52:53 2016 |
Steve | Update | PEM | leaky roof |
Johannes found dripping water at the vac rack. It is safe. It is not catching anything. Actual precipitation was only 0.62" |
Attachment 1: leakyRoof.jpg
|
|
Attachment 2: leakyRoofA.jpg
|
|
12030
|
Thu Mar 10 16:32:45 2016 |
rana | Update | endtable upgrade | Inventory check |
Its not a good idea to use green mounts with green lasers. Steve should be able to get another copy of the EY doubler mount made up if we really don't have another one sitting in the Manasa end table box which Koji mentioned. |
12029
|
Thu Mar 10 16:29:32 2016 |
gautam | Update | endtable upgrade | Inventory check |
I did a quick sweep of the lab to find out what hardware has already been acquired for the X-end table upgrade. The attached PDF is an inventory check in the spirit of this elog.
Some things we have to decide:
- Are we okay with using the old green coloured faraday mount for the IR faraday? I have in hand a piece identical to the one used at the Y-end for the green faraday, that is red in colour, so I guess we can switch this out.
- The way in which the doubling oven is currently mounted at the X-end is using some posts cobbled together. The Y-end looks to have a custom mount machined for it (see Attachment #2). Do we want to go ahead and get something like this done?
- I suppose it is okay to reuse all the old optics (mirrors, lenses, harmonic separators) and PDs? It may be that we need to order som extra mirrors/lenses/posts (this will become clear once I do the layout)
I have not gotten around to planning the layout or doing drawings. I will try and first work through a mode-matching solution to make sure we have all the required lenses. It may be that we need some 1" or 2" mirrors as well. The beam from the lightwave NPRO is quite elliptical, but we have a number of cylindrical lenses in hand already if we decide we want to use these, so I guess we don't have to worry about this...
This is quite a preliminary list, and I will add/update over the coming days as I do more detailed planning, but have I missed out anything obvious? |
Attachment 1: Inventory_check.pdf
|
|
Attachment 2: Doubler_comparison.pdf
|
|
12028
|
Thu Mar 10 03:03:11 2016 |
ericq | Update | LSC | DRFPMI Power stable, but no RF handoff |
[ericq, Gautam]
We worked on getting the DRFPMI back up and running, hoping the ALS performance was good enough.
We did succeed in bringing in enough of the AO path to stabilize arm powers > 100, but failed at the full RF DARM handoff.
REFL165 angle was adjusted to -86 to minimize PRCL in the Q signal.
The AS110 signals are mysteriously huger than they used to be. Whitening gain reduced to 15dB from 27dB. Old trigger thresholds are still fine.
The new AUX X laser has a different sign for the temperature-> frequency coupling, so our usual convention of "beatnote goes up when temp slider goes up" meant the ALSX input matrix elements had to change sign.
We think the POPDC PD (which I think is the POP2F PD) may be miscentered, since in PRMI configuration, its maximum does not coincide with the REFLDC minimum, and leaves a sizeable TEM10 lobe on the REFL camera. This was a pain. |
12027
|
Tue Mar 8 18:22:20 2016 |
rana | Update | Green Locking | Laser swap - some improvement |
Why is the transmission of X green so low? Perhaps you can phase lock the IR and then scan the X frequency, using the X arm as the analyzer. i.e. put a slow ramp into MC2 to pull the PSL frquency and thus the green frequency. You can record a movie of the scan using the framegrabber and record the green transmission peaks to see how big the mode match is exactly (which modes are so big) |
12026
|
Mon Mar 7 23:51:36 2016 |
gautam | Update | Green Locking | Laser swap - some improvement |
Quote: |
Next steps in recovering ALS and trying to lock again
- Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
- The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
- PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
- We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...
|
Attachment #1
Since I could not determine how many volts at the LO input of the pomona box input corresponds to how many volts at the laser PZT, I measured the transfer function between these points using the Agilent network analyzer. The measured TF suggests that for a function generator output of 2Vpp, we get approximately 75mrad of phase modulation, which compares reasonably well with the value of 120mrad reported here. I did not attempt to further increase the LO output signal to push this number closer to 120mrad, as with 2Vpp from the function generator we get +7dBm at the mixer, which is what it wants - so I wanted to avoid any attenuators etc...
Attachments #2 and #3
After ensuring that we have appreciable phase modulation, I set out to measure the PDH OLTFs and adjust the gain on the uPDH boxes accordingly. The X end gain is at 6.0, and the Y end gain is at 4.0. Before measuring the Y-end OLTF, I adjusted the steering mirrors to increase GTRY to ~0.45. GTRX remains a paltry 0.05... But the UGFs seem satisfactory..
Attachment #4
Finally, I took the ALS noise spectrum for the green beats. The beat note amplitudes on the network analyzer in the control room are still puny compared to what we had, -40dBm for Y and -45dBm for X. But the phase tracker Q values are ~1000 and ~3000 for X and Y respectively, which are pretty close to what these were if memory serves me right. There may still be some room for optimization of the PDH loop gains etc, and we could perhaps look at lowering the gain of the REFL PD at the X end? I also have yet to do the sweep for the 3 temperatures at which we can find a beatnote and park at the middle one...
These spectra suggest we could even possibly try locking? We are approximately a factor of 3 above the reference for X and on par with the reference for Y....
Unrelated to this work: I also realinged the PMC, PMC transmission is now 0.730V up from ~0.65V. |
Attachment 1: PomonaTF.pdf
|
|
Attachment 2: XPDH.pdf
|
|
Attachment 3: YPDH.pdf
|
|
Attachment 4: greenbeat_20160307.pdf
|
|
12025
|
Mon Mar 7 20:40:02 2016 |
ericq | Update | CDS | FB down again |
We went and looked at the monitor plugged into FB. All kinds of messages were being spammed to the screen (maybe RAM errors), and nothing could be done to interrupt. Sadly, a hard reboot of FB was neccesary.
Video of error messages: https://youtu.be/7rea_kokhPY
After the reboot, it just took a couple of model restarts to get the CDS screen happy. |
12024
|
Sun Mar 6 15:24:05 2016 |
gautam | Update | CDS | FB down again |
I came in to check the status of the nitrogen and noticed that the striptool panels in the control room were all blank.
- PMC was unlocked but I was able to relock it using the usual procedure
- FB seems to be down: I was unable to ssh into it (or any of the FEs for that matter). I checked the lights on the RAID array, they are all green. I am holding off on doing a hard reboot of FB in case there is some other debugging that can be done first
- None of the watchdogs were tripped, but judging by the green spots on the mirrors, all of them are moving quite a bit. I've shutdown the watchdogs on all the optics except the MC mirrors, but the ITMs and ETMs still seem to be moving quite a bit.
I am leaving things in this state for now. It is unclear why this should have happened, it doesn't seem like there was a power glitch? |
Attachment 1: 58.png
|
|
12023
|
Sat Mar 5 23:31:01 2016 |
gautam | Update | Green Locking | Laser swap - some updates |
I've been a little behind on my elogs so here is an update of the end laser situation.
IR beat for X-end recovered
- The issue was optimizing the alignment into the fiber at the end table.
- Using Fluke fiber illuminator helped in aligning IR pickoff into mount. Useful note: there is an unused fiber running between the X-end and the PSL table, by connecting these at the PSL table, I was able to monitor the coupled power while remaining at the X-end.
- Another major issue was that one of the steering mirrors (marked "Y1" in Attachment #1) was mounted with AR coated side facing the beam. This was fixed by simply rotating the post, the mirror was not removed from its mount. I can only assume that this mirror is in this kind of mount because of space constraints.
- The fiber has a collimating telescope attached to the end of it. In principle, this gives us more angular acceptance while coupling the beam into the fiber, but as I found out, the acceptance is still tiny (I don't have a number to quantify it). Furthermore, the Fluke visual fault locator revealed that the lens in the collimating telescope is not set up great - when re-doing the X end table, we should fix this situation so as to have a fairly large collimated beam coming out of the fiber when illuminated from the other end, this would make the mode matching much easier.
- Bottom line: we have ~1.2 mW of IR light incident on the coupler at the end table, and ~400uW of IR power at the PSL table => coupling efficiency is ~30%, not stellar, but sufficient for now I guess. After the various splitters etc, there is about 160uW of EX IR light and ~300uW of PSL IR light incident on the beat PD, and the beat amplitude is about -9dBm.
AM/PM characterization of newly installed Lightwave
- Having recovered the IR beat, I set out to do the PM characterization for the end laser.
- Attachment #2 shows the electrical setup. The IR beat was piped to the X-end via an existing long cable that runs between the vertex and the endtable. Not shown in the diagram, but I used a 20dB coupler to keep track of the beat frequency on the HP spectrum analyzer while doing this measurement.
- I restricted myself to the range between 100kHz and 500kHz to do the scan, because it takes quite a while to do the scan with fine resolution (IF bandwidth = 10Hz).
- To calibrate the magnitude response to rad/V, I divided the output of the network analyzer (converting dB to absolute magnitude first) by the amplitude of the signal seen on the monitoring oscilloscope while the PLL is unlocked. This number was 96mV/rad.
- To confirm that the error signal spectrum is indeed a good approximation of the "plant" transfer function (i.e that 100kHz >> UGF of loop transfer function of the PLL), I measured the loop TF of the PLL - Attachment #3 suggests a UGF of ~ 16kHz, which means the assumption is reasonable.
- Excitation amplitude was -25dBm (which gave reasonable SNR), and 3 averages were taken.
- The AM measurement was done using the same procedure as detailed here - the DC block was used. The DC level of the PD output was 2.72 V. The excitation amplitude was 0dBm.
- Attachment #4 shows the AM response, PM response and PM/AM ratio
- The peak in the PM/AM ratio at 256620 Hz is compelling because it is not too sharp (and so we can be reasonably confident we are at a good operating point) and the PM response of 23.83 rad/V is also acceptable.
- As a consistency check, the PM response of ~30rad/V at 100kHz => PZT actuator gain is ~3MHz/V, which is in the region we expect it to be...
Next steps in recovering ALS and trying to lock again
- Having set the PDH modulation frequency to 256.62kHz, I took the spectrum of ALS noise using the IR beat (i.e. by piping the IR beat signal through the electronics the green beats usually go through - 6dB and 10dB attenuators were placed immediately after the beat PDs for the X and Y arms respectively, to make the signal levels compatible with the electronics), Attachment #5 unfortunately suggests that the noise performance is still poor, and I suspect the situation will be similar using the green beat (though I have not measured this yet).
- The modulation depth could be sub-optimal for the X-end PDH, I have to measure this and check that it is at an acceptable level. This will also tell me if I need to change the sum+HPF pomona box used to send the PDH control signal + piezo dither signal to the laser PZT. In order to do this, I need to know what the input impedance to the FAST control BNC is - the manual isn't very helpful, it just says the piezo has a capacitance less than 10,000pF. I suppose I will have to actually measure this.
- PDH loop OLTFs have to be re-measured for both ends to check that the servo gain's are appropriately placed.
- We know that the mode-matching into the arm for the X end is poor (I have yet to quantify this) - I suspect that the beam ellipticity is the main culprit. However, the DC transmitted power levels at the PSL table are comparable to (even slightly better than) the Y arm numbers, and so this cannot be the sole reason why the X-arm ALS noise is so much worse... I will continue my investigations next week...
|
Attachment 1: AUXxTelescope.png.png
|
|
Attachment 2: PM_setup.pdf
|
|
Attachment 3: PLLolg.pdf
|
|
Attachment 4: AMPM20160303.pdf
|
|
Attachment 5: IRbeat_20160304.pdf
|
|
12022
|
Sat Mar 5 10:37:48 2016 |
rana | Update | IOO | PSL Laser Opened |
Sorry, that was me; taking some photos of the PSL and EX mirrors.
Quote: |
PSL Table doors were open, and the laser shutter was closed.
Doors have been closed, laser has been opened.
|
|
12021
|
Fri Mar 4 13:52:41 2016 |
ericq | Update | IOO | PSL Laser Opened |
PSL Table doors were open, and the laser shutter was closed.
Doors have been closed, laser has been opened. |
12020
|
Fri Mar 4 11:07:10 2016 |
Steve | Update | safety | crane inspection completes 2016 safey audit |
All 3 cranes inspected by professional and load tested with 450 lbs at max reach.
|
Attachment 1: craneInsp2016.jpg
|
|
12019
|
Fri Mar 4 01:11:41 2016 |
gautam | Update | Green Locking | Laser swap - both green beatnotes found |
The good news: both green beatnotes have now been found. The problem was alignment on the green beat PD on the PSL table which I fixed. They are about -40dBm in amplitude (compare to -25dBm we used to see). But looking at the phase tracker Q output seems to suggest that there is adequate signal...
The bad news: the ALS noise still looks bad (see attachment)- I think the IR beat for the Y was perhaps marginally better. The beat amplitude for the X beat was optimized on the PSL table with the help of the oscilloscope. There may be some headroom for improvement with the Y beat.
I also did the AM/PM measurement for the replaced lightwave, chose an LO frequency based on this, and took the loop OLTF, plots to follow...
To do:
- Check Y-end PDH loop OLTF
- Optimize beat note amplitude of Y beat
- Align Y-green better to the arm using steering mirrors on the endtable.
- Double check calibration of PM/AM measurement and that I've picked the correct LO frequency/ I don't have any other ideas for improving the situation with the X beat though
|
Attachment 1: IR_beat_20160303_green.pdf
|
|
12018
|
Thu Mar 3 10:19:20 2016 |
Steve | Update | safety | safety audit 2016 |
Bob cleaned the safety glasses. They were sonicated in warm 2% Liquinox water for 10 minutes. Steve checked them by transmission measurement of 1064 nm at 150 mW
|
Attachment 1: checkedSG.jpg
|
|
12017
|
Thu Mar 3 01:25:50 2016 |
gautam | Update | Green Locking | Laser swap - 2 IR + 1 green beatnotes found |
[ericq, gautam]
Summary of work done tonight:
- The PDH setup at the Y-end has been restored after I had pulled the whole thing apart some weeks ago to see that nothing was obviously wrong with the uPDH box
- Adjusted the temperature of the Y-end laser such that a beatnote was obtained - I did this using the IR beat (the end laser temp wasn't updated after the PSL temp was changed recently)
- The Y green beatnote was found easily, there was no alignment on the PSL table necessary, though there is room to improve this situation (beatnote amplitude was ~ -35dBm though we are used to more like -25dBm)
- The X green beat remains elusive - I played around with the alignment onto the green beat PD at the PSL table for some time, and the two beams are aligned as far as I can tell given the constrained area available in that area. It may be that I have to clear some optics, do a rigorous near-field/far-field alignment of the two beams and then try again
- Since we had two strong (-5dBm for Y, -9dBm for X) IR beatnotes, we decided to take the ALS noise spectra for these. So as to not overload the amplifiers, a -10dB attenuator (-6dB) was placed directlty after the Y (X) IR beat PDs, before routing these signals through the usual green beat signal chain. Attached is the measured spectrum. The new values of the temperature sliders at which beatnotes can be found are : 1700 for X and -5990 for Y (spectra taken at these values).
To do:
- For both ends, find the three temperatures at which we have beatnotes, and choose the middle one
- PM characterization of AUX X laser - it may be that the excess noise in the X spectrum is due to sub-optimal PDH
- Align the Y green better at the endtable, also take an OLTF measurement for the Y PDH loop
- Re-check the alignment onto the green beat PD for the X beat
Remarks:
- The Lightwave laser controllers differ from the Innolight ones in that it is not possible to directly set the signal to the SLOW control BNC to 0, and have that as the new reference point. Rather, there seems to be some setpoint which is saved as a reference, and the moment any signal is applied to the SLOW control BNC, it adjusts the actual temperature w.r.t. this saved setpoint. I believe it is possible to update this setpoint (it is also possible to update the calibration of the power readout, this is an additional issue at the X end), but since this wasn't critical, I've left it as is for the moment...
- The ALS nosie spectrum for the Y arm IR beat is surprisingly good!
|
Attachment 1: IR_beat_20160303_2.pdf
|
|
12016
|
Wed Mar 2 17:42:19 2016 |
gautam | Update | Green Locking | Laser swap - some progress |
[Koji, Johannes, gautam]
With Koji's and Johannes' help, I managed to resolve the coupling the pick-off IR beam into the fiber at the X end. I will put up a more detailed elog about how this was done - but in summary, we have about 31% coupling efficiency into the fiber, which isn't stellar, but I felt this was adequate to find a beatnote. Koji also pointed out that the collimation telescope attached to the fiber at the X-end is poorly mounted - this is something to fix when we swap endtables, but this was not addressed right now because if we were to adjust this, we would also have to adjust the mode matching into the fiber.
I then attempted to tune the temperature to find the IR beatnote. While doing so, I noticed some strange features of the controller - there are essentially two display modes relevant to laser crystal temperature, one which allows us to change the setpoint and one which is an actual readback of the temperature (this one can't be adjusted). While tuning the temperature, I noticed that the latter display ("LT") did not change in value. On a hunch, I disconnected the "SLOW" control BNC on the front panel, and voila, I was able to tune the setpoint and observe the measured temperature shift accordingly. I was thus able to find a reasonably strong IR beatnote (-9dBm) at T ~ 44.6 deg C (the beat PD was set to 0dB attenuation, i.e. high gain mode). However, the moment I reconnected the SLOW control BNC, the beatnote vanished (it gradually shifted out of range of the HP network analyzer), and the same thing happens if I terminate the SLOW control BNC connector! I don't understand this behaviour, as the manual says that the range of voltages accepted to this input is +/-10V, so I would assume 0V means do nothing, but clearly this isn't the case, as the beatnote is being shifted in frequency by > 1GHz, and the tuning coefficient is listed as 5GHz/V in the manual. This situation needs further investigation.
Since I had a reasonable IR beatnote setup, I returned the HP analyzer to the control room and tried to see if a green beatnote was present as well - I first ran ASS, then maximized the green transmission using the PZT mirrors, but no beatnote is evident. The contrast isn't great, the ratio of AUX power to PSL power on the green beat PD is something like 5:1, so this probably requires some tuning as well. I will update this elog after today evening's activities... |
12015
|
Wed Mar 2 10:09:28 2016 |
Steve | Update | safety | professional crane inspection |
The crane inspection is scheduled for this coming Friday from 8-12
|
12014
|
Tue Mar 1 09:37:15 2016 |
steve | Update | VAC | vac rack UPS batteries replaced |
Amstron batteries replaced after 11 months with SP-12-5.5HR, 2 years warranty from replaceUPSbattery.com
Quote: |
Batteries replaced after 3.5 years with Amstron AP-1250F2, 8x 12V 6Ah
Quote: |
APC Smart -UPS 2200 model: SUA2200RM2U batteries were replaced by compatible RBC43, 8x 12V5A
|
|
Note: the replace battery LED did not go out ( well pasted 24 hrs ) till the self test bottom was hold down for 2-3 sec |
Attachment 1: vacRackBat.jpg
|
|
12013
|
Mon Feb 29 17:17:26 2016 |
gautam | Update | Green Locking | Laser swap - still no green beatnote |
I continued the hunt for a green beatnote today - I decided to take the output from the RF amplifiers sitting on the PSL table and directly connect it to the analyzer in the control room while I swept the temperature of the end laser 10,000 counts on either side of a temperature at which I had taken this measurement - so I expect the beatnote should be found somewhere in this neighbourhood. But I did not see any peaks throughout the sweep. I re-checked that the mode overlap onto the BBPD is reasonable. We have considerably less transmitted green power from the arm now than we did before the laser swap (by a factor of ~3) but I still expected to see some sort of beat signal.
It would be handy to have the IR beat set up as well for this process, but as mentioned in a previous elog, I was getting only ~0.1 mW of IR power incident on the coupler at the end table last week. As I had suspected, tweaking the alignment of the steering optics for the pick-off IR beam after the doubler improved the situation somewhat, and I am now getting about 1mW of IR power incident on the coupler at the end table. But I've not been able to adjust the alignment into the fiber at the end such that I get any IR light at the PSL table. |
12012
|
Fri Feb 26 01:52:44 2016 |
gautam | Update | Green Locking | Laser swap - Green PDH OLTF |
I spent some more time today trying to optimize the modulation frequency and amplitude for the X end PDH, and the alignment/mode-matching of the green to the arm. Some notes:
- After my best efforts to tweak the alignment and mode-matching into the arm by using the two lenses on translational stages, I was able to get the green TRX up to about 0.06. As mentioned in a previous elog, this is much lower than what we had with the old setup, even though we have more green power going into the arm now. However, the mode looks pretty bright and clean on the monitors. Could the large ellipticity in the beam is the limiting factor now?
- I measured the transfer function (attachment #1) of the PDH loop once I had settled on a modulation frequency and amplitude that I judged to be optimal (indicated on the plot). The UGF is ~7kHz. The PDH error signal as viewed on the oscilloscope is comparable to what we had with the Innolight. All this optimization was done empirically, I have yet to do the PM measurement. I can't seem to get more than 0.2 mW of IR arriving at the fiber coupler, the number I found in some older elogs is 2mW with the old setup.
- I did some alignment of the PSL green and the X arm green onto the beat PD on the PSL table. After the power glitches, the doubling ovens do not automatically turn on, I had turned on the end ovens earlier, and today I turned on the PSL oven. I noticed some strange behaviour initially - though the setpoint was 36.9 deg C, when I enabled the heater, there was a large overshoot (it went to almost 50deg C). I disabled the heating at this point, and re-enabled it once the oven had cooled down to ~35 deg C. I didn't observe anything like this while turning on the end ovens. But the PID parameters at the PSL table are very different, so perhaps this large overshoot and ringing is to be expected. In any case, I managed to get this working. But I was not able to find a beatnote tonight.
To do:
- Verify that the two beams are aligned on the beat PD - I think I've done this carefully by checking the near and far-field, but I will double check.
- Find the beat note and look at the ALS noise performance with this new setup to see if it is usable even though GTRX is only 20% of what it used to be..
- Fix the coupling of the IR pickoff into the fiber at the endtable. Once this is done, I can do the PM measurement, and finding a beatnote may be easier given the IR beat PDs have a much wider bandwidth...
|
Attachment 1: X_PDH_OLTF_20160225.pdf
|
|
12011
|
Thu Feb 25 11:32:04 2016 |
gautam | Update | General | Power Glitch again |
Quote: |
10:15 power glitch today. ETMX Lightwave and air conditions turned back on
|
The CDS situation was not as catastrophic as the last time, it was sufficient for me to ssh into all the frontends and restart all the models. I also checked that monit was running on all the FEs and that there was no date/time issues like we saw last week. Everything looks to be back to normal now, except that the ntpd process being monitored on c1iscex says "execution failed". I tried restarting the process a couple of times, but each time it returns the same status after a few minutes. |
12010
|
Thu Feb 25 11:02:36 2016 |
Steve | Update | General | Power Glitch again |
Quote: |
Chiara reports an uptime of >195 days, so its UPS is working fine 
FB, megatron, optimus booted via front panel button.
Jetstor RAID array (where the frames live) was beeping, since its UPS failed as well. The beep was silenced by clicking on "View Events/Mute Beeper" at 192.168.113.119 in a browser on a martian computer. I've started a data consistency check via the web interface, as well. According to the log, this was last done in July 2015, and took ~19 hrs.
Frontends powered up; models don't start automatically at boot anymore, so I ran rtcds start all on each of them.
All frontends except c1ioo had a very wrong datetime, so I ran sudo ntpdate -b -s -u pool.ntp.org on all of them, and restarted the models (just updating the time isn't enough). There is an /etc/ntp.conf in the frontend filesystem that points to nodus, which is set up as an NTP server, but I guess this isn't working.
PMC locking was hindered by sticky sliders. I burtrestored the c1psl.snap from Friday, and the PMC locked up fine. (One may be fooled by the unchanged HV mon when moving the offset slider into thinking the HV KEPCO power supplies need to be brought down and up again, but it's just the sliders)
Mode cleaner manually locked and somewhat aligned. Based on my memory of PMC camera/transmission, the pointing changed; the WFS need a round of MC alignment and WFS offset setting, but the current state is fine for operation without all that.
|
10:15 power glitch today. ETMX Lightwave and air conditions turned back on |
Attachment 1: powerGlitch.png
|
|
12009
|
Wed Feb 24 19:29:13 2016 |
gautam | Update | Green Locking | Laser swap - Green PDH locked |
After the discussion at the meeting today, I decided to try and lock the green by sweeping through PZT dither frequencies in the vicinity of 200kHz without worrying about the AM/PM ratio for now. I was able to lock the PDH loop relatively quickly, at an empirically determined PZT dither frequency of 213.873kHz, 2Vpp (the amplitude was copied from the value at the Y-end). For today's efforts, I borrowed the sum+HPF pomona box from the Y-end, I will make a replica given that we are using Lightwave lasers at both ends now. After adjusting the PZT sliders and lenses on the translational stages at the endtable to maximize the green transmission as best as I could, I was able to get GTRX up to about 0.07 - this is far off from the value of ~0.25-0.3 I seem to remember us having with the old setup, even though we have more green light into the arm cavity. I will take a measurement of the loop transfer function to see what sort of bandwidth we have... |
12008
|
Wed Feb 24 10:27:23 2016 |
Steve | Update | safety | safety audit 2016 |
Safety audit went smothly.
Crane inspection is scheduled for March 4
Safety glasses will be measured before April 1
|
Attachment 1: audit2016.jpg
|
|
12007
|
Wed Feb 24 09:06:25 2016 |
Steve | Update | Treasure | the way it happened |
Celebration
|
Attachment 1: IMG_0120.JPG
|
|
12006
|
Tue Feb 23 23:01:16 2016 |
gautam | Update | Green Locking | Laser swap - Green PDH locking |
Given that we were seeing green flashes in the arms, I tried to see if I could get the green locked to the arm in a nice mode. For a start, I tried hooking up the PDH box and LO using the same settings as was being used previously. However, this did not work. I suppose we will have to do the whole AM/PM measurement for the Lightwave as well before we can determine what would be a suitable frequency for the LO. The AM measurement was relatively straightforward, I just repeated the same steps as detailed here. The two attachments show the AM response (one from 10kHz to 5MHz, the other for a narrower range of 100kHz to 1MHz, both with an excitation amplitude of 0dBm). To see if I could guess some sweetspot for operation, I tried setting the LO frequency to the two marked notch frequencies but was unsuccessful in getting the PDH lock going. At the moment, the alignment for the optics that picks off the IR after the doubler and routes it to the fiber are ccompletely misaligned, I will align these and do the PM measurement tomorrow and then we should conclusively be able to say what the appropriate frequency is to actuate on the PZT.
Unrelated to this work: the KEPCO high voltage power supply that drives the green steering mirror PZTs was switched off - I suppose this has been the case since the power outage last week. I turned it back on and reset it to the nominal settings: Vout = 100V, and Imax_out = 10mA, the driver board is currently drawing ~7mA which I judged to be consistent with the values labelled on the unit. |
Attachment 1: AM_scan.pdf
|
|
Attachment 2: AM_scan_zoomed.pdf
|
|
12005
|
Tue Feb 23 17:46:09 2016 |
gautam | Update | SUS | EQ |
Looks like another EQ 4.8 took out all the watchdogs, I've restored them, everything looks alright and doesn't look like any magnets got stuck this time... |
Attachment 1: SUS_Summary_23Feb.png
|
|
12004
|
Tue Feb 23 09:51:34 2016 |
Steve | Update | SUS | ITMX free |
Good job Q! Give us more details please or play Rachmaninoff Piano Concernto #2
Atm2, EQ 3.5 and 4.1
confirming ITMX is really free
Quote: |
Taking inspiration from J. Lewis et. al, ITMX has been freed.
Even better, I'm now seeing Green flashes in the X arm. 
|
|
Attachment 1: ITMX_FREED.png
|
|
Attachment 2: ITMX_FREE.png
|
|
12003
|
Mon Feb 22 23:00:21 2016 |
ericq | Update | SUS | ITMX free, green is flashing |
Taking inspiration from J. Lewis et. al, ITMX has been freed.
Even better, I'm now seeing Green flashes in the X arm.  |